# Fermionic spinon theory of the hourglass spin excitation spectrum of the cuprates

### v1

@@ 3 Discussion
In this section, we analyze the spin-structure factor obtained from our model and compare it with experimental observations in various cuprate compounds. We assume that the dominant contribution to the spin-structure factor comes from the third layer which is a $\pi$-flux spin liquid described by the Hamiltonian in Eq. (3). The system is treated within the mean-field approximation, neglecting the SU(2) fluctuations of the gauge field. Fig. 1 shows the imaginary part of the spin susceptibility for the third layer. A similar calculation was previously done in Ref. [Ma2018]. As shown, there is a gapless spinon continuum centered at $(\pi,\pi)$ in momentum space, exhibiting a Dirac-like dispersion.
@@
We examine the effect of the interacting part of the Hamiltonian using the results of the previous section, see Eq. (13). Incorporating interactions within the RPA framework leads to a renormalized spin susceptibility given by Eq. (14). Fig. 3 demonstrates that the including interactions leads to an emergent collective excitation (a triplon branch) below the spinon continuum. We choose Heisenberg couplings $J=0.235,J^{\prime}=-0.3J,J^{\prime\prime}=0.6J$$J^{\prime\prime\prime}=0.02J$, $J^{\prime\prime\prime\prime}=0.45J$ to achieve the dispersion that reproduces the characteristic hourglass-shaped spectrum. This type of spectrum has been observed in numerous neutron-scattering experiments on cuprate materials[Hayden2004, Vignolle2007, Chan2016].
@@
Importantly, the hourglass profile arises from averaging the CDW order over both spatial directions. Without this averaging, two distinct triplon branches appear, as illustrated in Fig. 4. Panels (a) and (b) correspond to the CDW order along the $x$ and $y$ directions correspondingly. In panel (a), the triplon branch exhibits a minimum at $(\pi,\pi)$, whereas in panel (b) the minimum is shifted away from $(\pi,\pi)$. The combination of the two produces the characteristic hourglass-shaped dispersion.
@@
Motivated by neutron-scattering experiments in a magnetic field [Coldea1997], we calculate the spin susceptibility in our model in the presence of a finite magnetic field in Appendix C. We find that the hourglass-like triplon branch splits into two branches, which are shifted in opposite directions by an energy of $2\Delta E_{z}$ due to Zeeman coupling. Although experiments require strong magnetic fields to resolve this splitting, such conditions may become possible in the future.
Finally, we comment on an alternative theoretical approach capable of reproducing the hourglass-shaped magnetic response, discussed in Appendix D. In particular, we employ the phenomenological theory of coupled spin and charge fluctuations introduced in Ref. [Vojta2006], to obtain an hourglass-shaped triplon branch outside the spinon continuum. Although this approach is valuable for its simplicity, the RPA-based susceptibility calculated in the main text offers a more direct and microscopic route to reproduce the experimental observations.
@@ 4 Conclusions
Near the charge-ordering wavevector, a RPA analysis of the interaction between the spinons produces the ‘hourglass’ spectrum observed by neutron scattering [Hayden2004, Vignolle2007, Chan2016]. Above the hourglass we obtain additional scattering from the spinon continuum which we hope will be studied in subsequent experiments.
The spinons are also visible in at other momenta, both in the form of RPA bound states and two-particle continua. This spectrum is a plausible explanation [SSZaanen] for the high energy scattering observed by RIXS experiments [Keimer11, Hayden19, Keimer25], and is present at similar energies. But the detailed form of our high energy spectrum appeares different from current observations e.g. the high intensity scattering disperses downwards from $(\pi,0)$ to $(0,0)$ in the RIXS observations [Hayden19, Keimer11], while that in Fig. 3 disperses upwards. We believe this is a consequence of the rather simple condensed form of the Higgs field, $B$, in a period 4 CDW state in our computation. It will be of interest to extend our study to the case of fluctuating CDW states using $B$ and SU(2) gauge fields, along the lines of Ref. [Sayantan25].
@@
  • The FL* theory accounts for the dispersion of the gapped electronic excitations near momentum $(\pi,0)$ (the ‘antinode’), with the spinons contributing to the broad linewidth [Mascot2022].
@@
  • The existence of the FL* state allows the possibility of an underlying quantum phase transition to a Fermi liquid (FL) without any symmetry-breaking order parameter [FLS1, Zhang2020]. Both FL* and FL have instabilities to the same $d$-wave superconductor, and so there is no quantum criticality within the superconducting state. But a strange metal state appears above the superconducting $T_{c}$ from FL*-FL quantum criticality in the presence of random interactions [Patel2]: such a state matches transport [Li:2024kxr], thermopower [LPS24], and Hall angle [Davide25] observations.
@@
  • The nodal quasiparticles of the low temperature $d$-wave superconducting state have strongly anisotropic velocities, with $v_{F}\gg v_{\Delta}$, where $v_{F}$ is the velocity along the zone diagonal, while $v_{\Delta}$ is the velocity along the orthogonal direction. This is difficult to understand from a $d$-wave superconductor descending from a spin liquid with massless Dirac fermion spinons [Zhang88, KotliarLiu88, IvanovSenthil, LeeWenRMP], as the spinons have nearly isotropic velocities. However, anisotropic nodal quasiparticles are obtained naturally by confining the FL* state with such a spin liquid background [Chatterjee16, Christos2024]: the spinons annihilate the Bogoliubov quasiparticles on the ‘backsides’ of the pockets, and observed Bogoliubov quasiparticles of the $d$-wave superconductor are those on the ‘front side’ of the pockets.
@@
  • The hole-doped quasiparticles display quantum oscillations at low temperatures and high magnetic fields. These oscillations are associated with electron pockets, as indicated by the negative Hall co-efficient in this regime. These oscillations are compatible with a model of electron pockets induced by bi-directional charge density wave order [HS11] only after accounting for the influence of spinons [BCS24].
@@
  • The observed structure of the vortex core in the underdoped cuprates is clearly distinct from that of the BCS theory of a $d$-wave superconductor emerging from a Fermi liquid [WM95]. The key differences can be described by a theory of the BCS $d$-wave superconductor emerging from a confinement transition of the FL* state [ZhangSS24, Sayantan25].
@@
  • The hole pockets survive in a $T=0$ FL* state with a non-zero quasiparticle residue around the hole pocket, even in the presence of the coupling to spinons [Mascot2022]. Recent work [Sayantan25] examined thermal fluctuations of a SU(2) gauge theory describing the spinons, at temperatures above a superconducting or charge-ordered ground state (the spinons are confined in these low temperature states). They showed that such thermal fluctuations do indeed convert the photoemission spectral weight to that of the observed ‘Fermi arcs’ in the intermediate temperature pseudogap regime.
@@
  • Finally, we note the evidence from ADMR and the Yamaji effect [fang_admr_2022, chan_yamaji_2024, Zhao_Yamaji_25], which was discussed in Section 1.
@@ Appendix D Alternative approach
In the main part of the paper we obtained the hourglass-like triplon branch as a collective mode of the RPA susceptibility. In this Appendix, we demonstrate that similar features exist in a model of fluctuating stripe charge order, proposed in [Vojta2006]. We use a phenomenological Lagrangian to describe the coupling between spin and charge fluctuations, which takes the following form:

### v2

@@ 3 Discussion
In this section, we analyze the spin-structure factor obtained from our model and compare it with experimental observations in various cuprate compounds. We assume that the dominant contribution to the spin-structure factor comes from the third layer which is a $\pi$-flux spin liquid described by the Hamiltonian in Eq. (3). The system is treated within the mean-field approximation, neglecting the SU(2) fluctuations of the gauge field. Fig. 1 shows the imaginary part of the spin susceptibility for the third layer. A similar calculation was previously done in Ref. [4]. As shown, there is a gapless spinon continuum centered at $(\pi,\pi)$ in momentum space, exhibiting a Dirac-like dispersion.
@@
We examine the effect of the interacting part of the Hamiltonian using the results of the previous section, see Eq. (13). Incorporating interactions within the RPA framework leads to a renormalized spin susceptibility given by Eq. (16). Fig. 3 demonstrates that the including interactions leads to an emergent collective excitation (a triplon branch) below the spinon continuum. We choose Heisenberg couplings $J=0.235,J^{\prime}=-0.3J,J^{\prime\prime}=0.6J$$J^{\prime\prime\prime}=0.02J$, $J^{\prime\prime\prime\prime}=0.45J$ to achieve the dispersion that reproduces the characteristic hourglass-shaped spectrum. We note that at half-filling, where the system is a Mott insulator, the coupling $J$ is exactly the nearest-neighbor antiferromagnetic Heisenberg exchange coupling, which has been estimated to be about 200 meV. This type of spectrum has been observed in numerous neutron-scattering experiments on cuprate materials[34, 35, 36].
@@
Importantly, the hourglass profile arises from averaging the CDW order over both spatial directions. The motivation for this is the following: the correlation length in most cuprate materials is quite small($100\text{\,}\mathrm{\SIUnitSymbolAngstrom}$\$), and in any real sample there are domains with various order orientation [55, 56]. A similar averaging procedure was done in Ref. [48] to study the hourglass shaped spectrum of LBCO. Without this averaging, two distinct triplon branches appear, as illustrated in Fig. 4. Panels (a) and (b) correspond to the CDW order along the $x$ and $y$ directions correspondingly. In panel (a), the triplon branch exhibits a minimum at $(\pi,\pi)$, whereas in panel (b) the minimum is shifted away from $(\pi,\pi)$. The combination of the two produces the characteristic hourglass-shaped dispersion.
@@
Motivated by neutron-scattering experiments in a magnetic field [60], we calculate the spin susceptibility in our model in the presence of a finite magnetic field in Appendix C. We find that the hourglass-like triplon branch splits into two branches, which are shifted in opposite directions by an energy of $2\Delta E_{z}$ due to Zeeman coupling. Although experiments require strong magnetic fields to resolve this splitting, such conditions may become possible in the future.
Finally, we comment on an alternative theoretical approach capable of reproducing the hourglass-shaped magnetic response, discussed in Appendix D. In particular, we employ the phenomenological theory of coupled spin and charge fluctuations introduced in Ref. [49], to obtain an hourglass-shaped triplon branch outside the spinon continuum. Although this approach is valuable for its simplicity, the RPA-based susceptibility calculated in the main text offers a more direct and microscopic route to reproduce the experimental observations.
@@ 4 Conclusions
Near the charge-ordering wavevector, a RPA analysis of the interaction between the spinons produces the ‘hourglass’ spectrum observed by neutron scattering [34, 35, 36]. Above the hourglass we obtain additional scattering from the spinon continuum which we hope will be studied in subsequent experiments.
The spinons are also visible in at other momenta, both in the form of RPA bound states and two-particle continua. This spectrum is a plausible explanation [47] for the high energy scattering observed by RIXS experiments [44, 45, 46], and is present at similar energies. But the detailed form of our high energy spectrum appeares different from current observations e.g. the high intensity scattering disperses downwards from $(\pi,0)$ to $(0,0)$ in the RIXS observations [45, 44], while that in Fig. 3 disperses upwards. We believe this is a consequence of the rather simple condensed form of the Higgs field, $B$, in a period 4 CDW state in our computation. It will be of interest to extend our study to the case of fluctuating CDW states using $B$ and SU(2) gauge fields, along the lines of Ref. [11].
@@
  • The FL* theory accounts for the dispersion of the gapped electronic excitations near momentum $(\pi,0)$ (the ‘antinode’), with the spinons contributing to the broad linewidth [7].
@@
  • The existence of the FL* state allows the possibility of an underlying quantum phase transition to a Fermi liquid (FL) without any symmetry-breaking order parameter [14, 28]. Both FL* and FL have instabilities to the same $d$-wave superconductor, and so there is no quantum criticality within the superconducting state. But a strange metal state appears above the superconducting $T_{c}$ from FL*-FL quantum criticality in the presence of random interactions [61]: such a state matches transport [62], thermopower [63], and Hall angle [64] observations.
@@
  • The nodal quasiparticles of the low temperature $d$-wave superconducting state have strongly anisotropic velocities, with $v_{F}\gg v_{\Delta}$, where $v_{F}$ is the velocity along the zone diagonal, while $v_{\Delta}$ is the velocity along the orthogonal direction. This is difficult to understand from a $d$-wave superconductor descending from a spin liquid with massless Dirac fermion spinons [65, 66, 67, 68], as the spinons have nearly isotropic velocities. However, anisotropic nodal quasiparticles are obtained naturally by confining the FL* state with such a spin liquid background [69, 9]: the spinons annihilate the Bogoliubov quasiparticles on the ‘backsides’ of the pockets, and observed Bogoliubov quasiparticles of the $d$-wave superconductor are those on the ‘front side’ of the pockets.
@@
  • The hole-doped quasiparticles display quantum oscillations at low temperatures and high magnetic fields. These oscillations are associated with electron pockets, as indicated by the negative Hall co-efficient in this regime. These oscillations are compatible with a model of electron pockets induced by bi-directional charge density wave order [70] only after accounting for the influence of spinons [10].
@@
  • The observed structure of the vortex core in the underdoped cuprates is clearly distinct from that of the BCS theory of a $d$-wave superconductor emerging from a Fermi liquid [71]. The key differences can be described by a theory of the BCS $d$-wave superconductor emerging from a confinement transition of the FL* state [72, 11].
@@
  • The hole pockets survive in a $T=0$ FL* state with a non-zero quasiparticle residue around the hole pocket, even in the presence of the coupling to spinons [7]. Recent work [11] examined thermal fluctuations of a SU(2) gauge theory describing the spinons, at temperatures above a superconducting or charge-ordered ground state (the spinons are confined in these low temperature states). They showed that such thermal fluctuations do indeed convert the photoemission spectral weight to that of the observed ‘Fermi arcs’ in the intermediate temperature pseudogap regime.
@@
  • Finally, we note the evidence from ADMR and the Yamaji effect [31, 32, 33], which was discussed in Section 1.
@@ Appendix D Alternative approach
In the main part of the paper we obtained the hourglass-like triplon branch as a collective mode of the RPA susceptibility. In this Appendix, we demonstrate that similar features exist in a model of fluctuating stripe charge order, proposed in [49]. We use a phenomenological Lagrangian to describe the coupling between spin and charge fluctuations, which takes the following form: